# DFT modelling of stacking faults in hexagonal and cubic GaN

### v1

@@ Abstract
We have performed density functional theory (DFT) calculations to characterize the energetics, and the atomic and electronic structure, of stacking faults in GaN, both in the stable hexagonal wurtzite (wz) phase and in the metastable cubic zincblende (zb) phase. In wz GaN, SFs on the (0001) planes can be divided into three different intrinsic stacking faults (I1, I2, and I3) and one extrinsic stacking fault (E). In zb GaN, SFs form along (111) directions, giving one type each of intrinsic, extrinsic and twin SFs. Based on the calculated formation energy, I1 is the most stable SF of wz GaN in agreement with experiment. For zb GaN, the intrinsic stacking fault is the most dominant planar defect. To characterize the effect of the stacking faults on the electronic structure of the material, we examined the band density. We found that the bands near the valence band maximum in wz GaN are localised on the Ga-polar side of the stacking fault (i.e. on the Ga side of the Ga-N bonds perpendicular to the SF), with the bands near the conduction band minimum more on the N-polar side, though somewhat delocalised. We found the opposite trend in zb GaN; this behaviour is caused by a redistribution of charge near the interface. We also show the band offsets for the stacking faults, finding that they are very sensitive to local conditions, but can all be described as type II interfaces, with the presence of a stacking fault reducing the gap locally.
@@ I Introduction
III-V nitrides are increasingly important in electronic and optoelectronic devices [12, 18]. These applications are affected by a number of material problems, mostly due to the lack of suitable substrates on which the materials are grown. Growth on mismatched substrates causes the epilayers to contain a very high concentration of extended defects, i.e., stacking faults, up to five orders of magnitude higher than in other materials used for optoelectronic devices [21]. Although most commercial GaN devices are based on the stable $c$-oriented wurtzite (wz) GaN, these materials present a strong piezoelectric effect in the [0001] direction [2]. As a result of this phenomenon, band bending occurs through the quantum confined Stark effect [5, 26].
@@
Through epitaxial growth of GaN films on cubic substrates, a metastable cubic zincblende (zb) structure can be stabilized [35, 18], which is not polar like the wz phase, reducing the issues for optoelectronic devices. Recent studies show that, due to the absence of polarization fields and the lower band gap compared to wz GaN, zb GaN shows higher efficiency in longer wavelength emission (green, amber, red) [6]. The stacking faults in zb GaN can be treated as wz insertions in a zb matrix [6] which modify the optical properties of cubic GaN films and propagate into active layers. In quantum wells and electron blocking layers, the segregation of alloying elements has been observed at stacking faults, leading to the formation of quantum wires and polarized emission [11]. The SFs have been shown to facilitate electronic confinement and induce spontaneous ordering in p-type systems [22, 30].
@@
Many questions have yet to be answered concerning the high resolution characterization of SFs’ densities, alloy segregation to SFs and their influence on the optoelectronic properties of GaN-based devices. The density of stacking faults in basal (0001) planes were measured with x-ray diffraction (XRD) [2]. High resolution transmission electron microscopy (TEM) helped to understand zb/wz heterostructures, which result from the spontaneous polarization of wz GaN [20]. Recently, a combination of high-resolution scanning transmission electron microscopy (STEM) and energy dispersive x-ray spectrometry was used to investigate the effects of alloy segregation around stacking faults in a zb GaN light-emitting structure [40]. A key question remaining is to identify the change in atomic spacing at the SFs, and whether this plays a significant role in segregation phenomenon [11]: here, atomistic modelling can complement experiments, and provide insights beyond experimental resolution. DFT modelling has previously focused on the energetics and electronic structure of basal stacking faults of wz GaN [32, 3, 4] with forcefield modelling considering energetics in zb GaN [1]. However, there have been no DFT studies on SFs in zb GaN, and a detailed examination of the effects of stacking faults on the local electronic structure and band edges has not been performed.
@@
In this paper, our aim is to study the atomic structure and energetics of SFs, alongside their electronic structure and differences to bulk material in both wz and zb phases of GaN, and to compare to recent high-resolution STEM imaging of those SFs [40, 39]. The rest of the paper is laid out as follows: in Sec. II, the computational approach is presented; the results for the structural and energetic properties are compared in Sec. III; the electronic structure of the SFs is discussed in Sec. IV; Sec. V gives a brief summary and concludes the paper.
@@ III Physical structure and energetics
Taking the optimized lattice parameters and atomic coordinates of bulk structures, a series of simulations cells was constructed containing stacking faults, with progressively increasing separation between the stacking faults to test the convergence. For wz GaN, the three intrinic stacking faults I1, I2, I3 in isolation would be formed as …AB/CB…, …AB/CA…and …AB/C/BA…respectively. When modelled with periodic boundary conditions some adjustment is required, and we need two stacking faults in each cell. For n layers of each stacking sequence, we then have /nAB/nCB/, /nAB/nCA/C/ and /nAB/C/nBA/C/ where the stacking faults are in the middle and at the edges of the cell, and an extra C layer has been added to both I2 and I3 to restore periodicity. For the extrinsic stacking fault, we have …AB/C/AB…in isolation, and /nAB/C/ in a periodic cell (in this case, we need only one stacking fault per simulation cell).
For zb GaN, in isolation, the intrinsic, extrinsic and twin stacking faults are formed as …ABC/BC/ABC…, …ABC/B/ABC… and …ABC/BAC/ABC… respectively. We can accomodate a single stacking fault in a periodic cell in all three cases, which for n layers we write as /nABC/BC/, /nABC/B/ and /nABC/BAC/ in each case. These models are illustrated in Figs. 1 and 2.
Figure 1: Stacking faults in wurtzite GaN. (a) Intrinsic 1 (I1) /AB/CBCB/AB/, (b) Intrinsic 2 (I2) /AB/CACA/C/AB/, (c) Intrinsic 3 (I3) /AB/C/BABA/C/AB/ and (d) Extrinsic /ABAB/C/ABAB/. Note that these are highly shortened cells for illustration only.
@@
In zb GaN, however, all stacking faults have a negative formation energy; this is not surprising, as SFs represent a small wz inclusion in the zb lattice, and wz is the more stable crystal. All of the SFs have almost the same energy (certainly the difference between them is within the accuracy limit of these calculations) though the experimental observation is that the intrinsic stacking fault (SF) is the most abundant planar defect in epitaxial layers [11].
All stacking faults have a very low formation energy, implying that these defects may form very easily in both wz and zb GaN, though they are much more likely to form in the zb phase.
@@ I E T
Accurate determination of the stacking fault thickness ($l$) is essential for calculating the band profile of stacking faults functioning as quantum wells (QWs) [6]. The calculated stacking-fault thickness $l$ for the optimized models of all the considered types of SFs are provided in Table 3. Here we use the inter-planar spacing as discussed in a recent experimental investigation based on combined high-resolution scanning transmission electron microscopy and energy dispersive x-ray spectrometry [11]. At 300 K, the $c$ lattice parameter of bulk wz GaN and the lattice parameter of bulk zb GaN are measured to be 5.186 Å and 4.506 Å, respectively, corresponding to inter-planar spacings of 2.593 Å and 2.602 Å in wz-GaN and zb-GaN [14, 13, 31]. Bulk zb-GaN exhibits a larger inter-planar spacing between close-packed planes compared to bulk wz-GaN, which may be attributed to stronger Coulombic interactions between third-nearest-neighbor Ga and N atoms in the wurtzite phase. Our calculated lattice parameters for the bulk wz GaN are $a$ = 3.235 Å and $c$ = 5.284 Å and for bulk zb GaN, it is $a$ = 4.578 Å. The resulting inter-planar spacing d(0001) in bulk wz-GaN is 2.642 Å and d(111) in bulk zb-GaN is 2.644 Å. These show good agreement with the experimental measurements[14, 13, 31] though perhaps the difference between the phases is a little small. SFs can be seen as small insertions into the crystal of the other crystal type, and their equilibrium inter-planar spacing may, thus, be similar to the inter-planar spacing in the non-defective materials. In Table 3, the calculated largest ($l_{max}$) and smallest ($l_{min}$) distances of the SFs are presented. Our calculated stacking faults thicknesses range from 2.639 to 2.643 Å for wz GaN and 2.643 to 2.644 Å for zb GaN, showing a very small variation; plots of the evolution of the layer spacing along the $c$ axis of the simulation cell are given in the supporting information. (The alternative approach to define SF thickness using multiples of the Ga-N bond length [4, 19] gives similar insight but with less detail, so we do not use it.)
@@ IV Electronic structure
In the first instance, we plot band resolved densities to visualise the changes around the Fermi level in real-space. In the usual way, we define the density for band $n$ as $\rho_{n}(\mathbf{r})=\sum_{\mathbf{k}}w_{\mathbf{k}}|\psi_{n\mathbf{k}}(\mathbf{r})|^{2}$ where $w_{\mathbf{k}}$ is the weight for each point in the Brillouin zone. The band densities show similar trends for all SFs in each material, so we show here the bands for the most common SF in wz (I1) in Fig. 3 and in zb (I) in Fig. 4. Band densities for all other SFs are given in the supporting information.
@@
Figure 3: Band resolved densities of Intrinsic-1 stacking fault of wz GaN, here shown with six layers between stacking faults (/3AB/6CB/3AB/). The bands below the Fermi energy are shown in panels (a-d), and the bands above the Fermi energy are shown in panels (e-h). The isosurfaces were plotted at a density of 0.0005 electrons per Bohr3. Note that there are two stacking faults in the simulation cell, 25% and 75% of the way along the cell.
@@
Figure 4: Band resolved densities of Intrinsic stacking fault for zb GaN along (111), here shown with eight layers between stacking faults (/4ABC/BC/4ABC/). The bands below the Fermi energy are shown in panels (a-d), and the bands above the Fermi energy are shown in panels (e-h). The isosurfaces were plotted at a density of 0.0005 electrons per Bohr3. There is one stacking fault, halfway along the simulation cell.
@@
It is important to understand the nature of the band alignment around the SFs (whether a type I or type II interface) as it can affect the trapping of electrons and holes and hence the optoelectronic properties[9]. We compare the band edges in simulation cells with stacking faults to cells of perfect bulk, ensuring that the deep-lying non-dispersive $d$-bands are aligned as explained in Sec. II. The band edges are shown in Fig. 6 for both wz and zb. It should be noted that, as is standard for most DFT calculations, our calculated band gaps 1.65 eV (bulk wz GaN) and 1.52 eV (bulk zb GaN) are underestimated compared to the experimental values (3.51 eV and 3.30 eV for bulk wz and zb GaN respectively[37]). This underestimate is a well-known feature of standard DFT functionals, and our calculations are in good agreement with other DFT values[4].
We see that in zb, for all SFs, these defects are type II interfaces, with both valence band maximum and conduction band minimum shifted higher than in the bulk. For wz the same is true, though for the I1 SF (the most common) the conduction band minimum is extremely close to the energy in the bulk; the behaviour is likely to be very sensitive to local fields and perturbations of the electronic structure, so that hints type I behaviour might be seen. This is in broad agreement with existing simulation results[19, 32, 23, 33], though we are not aware of systematic testing of the size of simulation cell in other theory work. We also see that the gap is reduced by $\sim$0.01 eV in the wz I1 SF and by $\sim$0.04 eV in zb I SF. Experimental CL and PL results show a similar narrowing of the gap[16] from SFs in wz samples, though the magnitude is larger, at $\sim$0.9 eV.
@@ V Conclusions
We have performed a systematic study of stacking faults in wurtzite and zincblende phases of GaN using periodic DFT calculations. We have investigated all possible types of stacking faults structures, their stability and electronic properties to understand the structure properties relationship. Such investigation is of central importance since stacking faults are very common defects in III-nitride materials. From the structural relaxation, we can observe that the stacking faults have a rather small influence on lattice parameters and atomic positions; the inter-planar distance varies very slightly around the stacking faults.
@@
Finally, a detailed analysis of the electronic structure shows that the band gap of stacking faults is smaller than that of the ideal wurtzite and zincblende phases, with a positive offset. This change in the band edges is equivalent to a type II interface in all cases, though the most common wz stacking fault is highly sensitive.
@@ Appendix A Basis sets
[…]

### v2

@@ Abstract
We have performed density functional theory (DFT) calculations to characterize the energetics, and the atomic and electronic structure, of stacking faults in GaN, both in the stable hexagonal wurtzite (wz) phase and in the metastable cubic zincblende (zb) phase. In wz GaN, SFs on the (0001) planes can be divided into three different intrinsic stacking faults (I1, I2, and I3) and one extrinsic stacking fault (E). Based on the calculated formation energy, I1 is the most stable SF of wz GaN in agreement with experiment. In zb GaN, SFs form along (111) directions, giving one type each of intrinsic, extrinsic and twin SFs. In our calculations, the three types of SFs have similar formation energy. To characterize the effect of the stacking faults on the electronic structure of the material, we examined the band density. We found that the bands near the valence band maximum in wz GaN are localised on the Ga-polar side of the stacking fault (i.e. on the Ga side of the Ga-N bonds perpendicular to the SF), with the bands near the conduction band minimum more on the N-polar side, though somewhat delocalised. We found the opposite trend in zb GaN; this behaviour is caused by a redistribution of charge near the interface. We also show the band offsets for the stacking faults, finding that they are very sensitive to local conditions, but can all be described as type II interfaces, with the presence of a stacking fault reducing the gap locally in both wz and zb.
@@ I Introduction
III-V nitrides are increasingly important in electronic and optoelectronic devices [13, 19]. These applications are affected by a number of material problems, mostly due to the lack of cost-effective suitable substrates on which the materials are grown. Growth on mismatched substrates causes the epilayers to contain a very high concentration of extended defects, principally dislocations in c-type wz GaN [21, 24] and stacking faults in non-polar oriented wz [17, 42]. Most commercial GaN devices are based on the stable $c$-oriented wurtzite (wz) GaN, which shows a spontaneous polarisation; moreover, in the strained heterostructures required for devices, piezoelectric polarisation is induced [35]. As a result of this phenomenon, band bending occurs through the quantum confined Stark effect (QCSE) [5, 27].
@@
Through epitaxial growth of GaN films on cubic substrates, a metastable cubic zincblende (zb) structure can be stabilized [36, 19], which is not polar like the wz phase, providing a potential improvement in issues associated with the QCSE that occur in wz optoelectronic devices. Recent studies show that, due to the absence of polarization fields and the lower band gap compared to wz GaN, zb GaN may show higher efficiency in longer wavelength emission (green, amber, red) [10]. The stacking faults in zb GaN are often considered as wz insertions in a zb matrix [6] which modify the optical properties of zb GaN films and propagate into active layers (though we show here that this model is not correct). In quantum wells and electron blocking layers, the segregation of alloying elements has been observed at stacking faults, leading to the formation of quantum wires and polarized emission [12].
@@
Many questions have yet to be answered concerning the high resolution characterization of SFs’ structures, alloy segregation to SFs and their influence on the optoelectronic properties of GaN-based devices. Recently, a combination of high-resolution scanning transmission electron microscopy (STEM) and energy dispersive x-ray spectrometry was used to investigate the effects of alloy segregation around stacking faults in a zb GaN light-emitting structure [12]. A key question remaining is to identify the change in atomic spacing at the SFs, and whether this plays a significant role in segregation phenomena, which has been suggested as a possible mechanism for the different behaviour of Al and In [12]: here, atomistic modelling can complement experiments, and provide insights beyond experimental resolution. DFT modelling has previously focused on the energetics and electronic structure of basal stacking faults of wz GaN [32, 3, 4] with forcefield modelling considering energetics in zb GaN [1]. However, there have been no DFT studies on SFs in zb GaN, and no examination of the effects of stacking faults on the local electronic structure and band edges in either wz or zb GaN.
@@
In this paper, our aim is to study the atomic structure and energetics of SFs, alongside their electronic structure and differences to bulk material in both wz and zb phases of GaN, and to compare to recent experimental findings on those SFs (e.g. wz GaN [2, 42, 25] and zb GaN [41, 40]). The rest of the paper is laid out as follows: in Sec. II, the computational approach is presented; the results for the structural and energetic properties are compared in Sec. III; the electronic structure of the SFs is discussed in Sec. IV; Sec. V gives a brief summary and concludes the paper.
@@ III Physical structure and energetics
Taking the optimized lattice parameters and atomic coordinates of bulk structures, a series of simulations cells was constructed containing stacking faults, with progressively increasing separation between the stacking faults to test the convergence. For wz GaN, the three intrinsic stacking faults I1, I2, I3 in isolation would be formed as …AB/CB…, …AB/CA…and …AB/C/BA…respectively. When modelled with periodic boundary conditions some adjustment is required, and we need two stacking faults in each cell. For n layers of each stacking sequence, we then have /nAB/nCB/, /nAB/nCA/C/ and /nAB/C/nBA/C/ where the stacking faults are in the middle and at the edges of the cell, and an extra C layer has been added to both I2 and I3 to restore periodicity. For the extrinsic stacking fault, we have …AB/C/AB…in isolation, and /nAB/C/ in a periodic cell (in this case, we need only one stacking fault per simulation cell).
For zb GaN, in isolation, the intrinsic, extrinsic and twin stacking faults are formed as …ABC/BC/ABC…, …ABC/B/ABC… and …ABC/BAC/ABC… respectively. We can accommodate a single stacking fault in a periodic cell in all three cases, which for n layers we write as /nABC/BC/, /nABC/B/ and /nABC/BAC/ in each case. These models are illustrated in Figs. 1 and 2.
Figure 1: Stacking faults in wurtzite GaN. (a) Intrinsic-1 (I1) /AB/CBCB/AB/, (b) Intrinsic-2 (I2) /AB/CACA/C/AB/, (c) Intrinsic-3 (I3) /AB/C/BABA/C/AB/ and (d) Extrinsic /ABAB/C/ABAB/. Note that these are highly shortened cells for illustration only.
@@
In zb GaN, however, all stacking faults have a negative formation energy; this is not surprising, as SFs represent a small wz inclusion in the zb lattice, and wz is the more stable crystal. All of the SFs have almost the same energy (certainly the difference between them is within the accuracy limit of these calculations) though the experimental observation is that the intrinsic stacking fault (SF) is the most abundant planar defect in epitaxial layers [12]. All stacking faults have a very low formation energy, implying that these defects may form very easily in both wz and zb GaN, though they are much more likely to form in the zb phase. We note that the rotation of the Ga-N bonds visible in ABF (annular bright field) STEM images of the intrinsic stacking fault[40] is reproduced in the calculated charge density of the modelling.
@@ I E T
Accurate determination of the stacking fault thickness ($l$) is essential for calculating the band profile of stacking faults functioning as quantum wells (QWs) [6]. The calculated stacking-fault thickness $l$ for the optimized models of all the considered types of SFs are provided in Table 3. Here we use the inter-planar spacing as a measure of stacking-fault thickness, which is consistent with a recent experimental investigation based on high-resolution scanning transmission electron microscopy [12]. At 300 K, the $c$ lattice parameter of bulk wz GaN and the lattice parameter of bulk zb GaN are measured to be 5.186 Å and 4.506 Å, respectively, corresponding to inter-planar spacings of 2.593 Å and 2.602 Å in wz-GaN and zb-GaN [15, 14, 31]. Bulk zb-GaN exhibits a larger inter-planar spacing between close-packed planes compared to bulk wz-GaN, which may be attributed to stronger Coulombic interactions between third-nearest-neighbor Ga and N atoms in the wurtzite phase. Our calculated lattice parameters for the bulk wz GaN are $a$ = 3.235 Å and $c$ = 5.284 Å and for bulk zb GaN, it is $a$ = 4.578 Å. The resulting inter-planar spacing d(0001) in bulk wz-GaN is 2.642 Å and d(111) in bulk zb-GaN is 2.644 Å. These show good agreement with the experimental measurements[15, 14, 31] though perhaps the difference between the phases is a little small. SFs can be seen as small insertions into the crystal of the other crystal type, and their equilibrium inter-planar spacing may, thus, be similar to the inter-planar spacing in the non-defective materials. In Table 3, the calculated largest ($l_{max}$) and smallest ($l_{min}$) layer spacings of the SFs are presented. Our calculated stacking faults layer spacings range from 2.639 to 2.643 Å for wz GaN and 2.643 to 2.644 Å for zb GaN, showing a very small variation; plots of the evolution of the layer spacing along the $c$ axis of the simulation cell are given in the supporting information. (The alternative approach to define SF thickness using multiples of the Ga-N bond length [4, 20] gives similar insight but with less detail, so we do not present it.)
@@ IV Electronic structure
In the first instance, we plot band-resolved charge densities to visualise the changes around the Fermi level in real-space. In the usual way, we define the density for band $n$ as $\rho_{n}(\mathbf{r})=\sum_{\mathbf{k}}w_{\mathbf{k}}|\psi_{n\mathbf{k}}(\mathbf{r})|^{2}$ where $w_{\mathbf{k}}$ is the weight for each point in the Brillouin zone; the resulting plots are a solid state equivalent of a molecular orbital plot. These band densities show similar trends for all SFs in each material, so we show here the bands for the most common SF in wz (I1) in Fig. 3 and in zb (I) in Fig. 4. Band densities for all other SFs are given in the supporting information.
@@
Figure 3: Band-resolved densities of Intrinsic-1 stacking fault of wz GaN, here shown with six layers between stacking faults (/3AB/6CB/3AB/). The bands below the Fermi energy are shown in panels (a-d), and the bands above the Fermi energy are shown in panels (e-h). The isosurfaces were plotted at a density of 0.0005 electrons per Bohr3. Note that there are two stacking faults in the simulation cell, 25% and 75% of the way along the cell.
@@
Figure 4: Band-resolved densities of Intrinsic stacking fault for zb GaN along (111), here shown with eight layers between stacking faults (/4ABC/BC/4ABC/). The bands below the Fermi energy are shown in panels (a-d), and the bands above the Fermi energy are shown in panels (e-h). The isosurfaces were plotted at a density of 0.0005 electrons per Bohr3. There is one stacking fault, halfway along the simulation cell.
@@
It is important to understand the nature of the band alignment around the SFs (whether a type I or type II interface) as it can affect the trapping of electrons and holes and hence the optoelectronic properties[9]. We compare the band edges in simulation cells with stacking faults to cells of perfect bulk, ensuring that the deep-lying non-dispersive $d$-bands are aligned as explained in Sec. II. The band edges are shown in Fig. 6 for both wz and zb. It should be noted that, as is standard for most DFT calculations, our calculated band gaps 1.65 eV (bulk wz GaN) and 1.52 eV (bulk zb GaN) are significantly underestimated compared to the experimental values (3.51 eV and 3.30 eV for bulk wz and zb GaN respectively[38]). This underestimate is a well-known feature of standard DFT functionals, and our calculations are in good agreement with other DFT values[4].
We see that, for all SFs in zb, these defects are type II interfaces, with both valence band maximum and conduction band minimum shifted higher than in the bulk. For wz the same is true, though for the I1 SF (the most common) the conduction band minimum is extremely close to the energy in the bulk; the behavior is likely to be very sensitive to local fields and perturbations of the electronic structure, so that hints of type I behavior might be seen. This is in broad agreement with existing simulation results[20, 32, 23, 33], though we are not aware of systematic testing of the size of simulation cell in other theory work. We also see that the gap is reduced by $\sim$0.01 eV in the wz I1 SF and by $\sim$0.04 eV in zb I SF. This result is in contradiction to simple models of SFs in zb, assumed to be wz insertions into the zb matrix, which would lead to an increase of the gap. This highlights the importance of detailed electronic structure modelling of realistic models of materials. Experimental CL and PL results show bands associated with SFs in wz samples which are at lower energies than the band edge[22, 17], with shifts of $\sim$0.06–0.18 eV.
@@ V Conclusions
We have performed a systematic study of stacking faults in wurtzite and zincblende phases of GaN using periodic DFT calculations. We have investigated all possible types of stacking faults structures, considering their stability and electronic properties to understand the structure properties relationship. This investigation is of central importance since stacking faults are very common defects in zb III-nitride materials, and non-polar oriented wz materials. From the structural relaxation, we can observe that the stacking faults have a rather small influence on lattice parameters and atomic positions; the inter-planar distance varies very slightly around the stacking faults.
@@
Finally, a detailed analysis of the electronic structure shows that, in contrast to standard models, the band gap of stacking faults is smaller than that of the ideal wurtzite and zincblende phases, with a positive offset. This change in the band edges is equivalent to a type II interface in all cases, though the most common wz stacking fault is highly sensitive to local conditions.
@@ Appendix A Basis sets
The pseudo atomic orbital (PAO) basis sets[8] used for Ga and N were double-zeta plus polarization (DZP) basis sets, generated with the equal radius setting in Conquest. The gallium pseudopotential includes the $3d$ electrons as semi-core states, with the radii of the PAOs in Bohr as follows: $3d$: 3.58 $a_{0}$; $4s$ and $4p$: 8.06 and 4.33 $a_{0}$; 4d (polarization): 8.06 $a_{0}$. For N the radii are: $2s$ and $2p$: 5.48 and 2.81 $a_{0}$; 3d (polarization): 5.48 $a_{0}$. For indium, the pseudopotential includes the $4d$ electrons as semi-core states, with the radii of the PAOs in Bohr as follows: $4d$: 4.11 $a_{0}$; $5s$ and $5p$: 8.41 and 4.63 $a_{0}$; 5d (polarization): 8.41 $a_{0}$. For aluminium, the radii of the PAOs in Bohr are as follows: $3s$ and $3p$: 8.24 and 4.48 $a_{0}$; 3d (polarization): 8.24 $a_{0}$.